4√4x-8 = 3
I don't understand .. I'm supposed to solve 4 x.. I need help.. thank you!


Answer by unlockmath(1688)

Hello,
For this 4√4x-8 = 3
we want to first add 8 to get:
4√4x = 11
Now divide by 4:
√4x = 11/4
Now square both sides to get:
4x=121/16
Divide by 4:
x=121/64
Make Sense?
RJ
www.math-unlock.com